# GAD-Generative Learning for HD Map-Free Autonomous Driving

### v1

@@ A.6 Lattice Sampler
A.7 Retrieval-based Sampler
The retrieval based sampler does not depend on the map or lane information to generate trajectories, instead, it retrieves a cluster of trajectories whose initial state are close to the SDV’s current state from a pre-established expert trajectory dataset.
@@ A.7 Retrieval-based Sampler
To subdivide the original trajectories into different bins, we use the initial velocity v, acceleration a and curvature $\kappa$, with respective bin sizes of $1.0\frac{m}{s}$, $0.5\frac{m}{s^{2}}$, $0.01\frac{1}{m}$. In each bins, trajectories are clustered into 2000 sets and the closest trajectory to each cluster center is retained. Finally, the trajectories from each subset will be aggregated to the comprehensive expert trajectory dataset.
@@
$\textit{D}=\beta_{1}\cdot|v-v_{SDV}|+\beta_{2}\cdot|a-a_{SDV}|+\beta_{3}\cdot|\kappa-\kappa_{SDV}|$ (1)
@@ A.8 Hardware
A.9 Task Perf
@@
A.10 Road Test
Our road tests are conducted in the Jiading District of Shanghai, China, known for ”Shanghai Automobile City”. The detailed testing route is illustrated in Figure 4, encompassing straight stretches, lane changes, intersections, and turning scenes involving various traffic participants.
The route begins at No.1688 Yecheng Road, a road with two lanes in each direction, encounters frequent cars and electric bicycles. It passes through two intersections controlled by traffic lights before entering a three-lane structure and making a careful lane change to the right-most lane, where sometimes occupied by static vans. It then queues before the stop line and then turns onto Shengxin Road, a four-lane-per-direction road heavily congested commuter route with significant traffic flows in all directions.
@@ A.10 Road Test
The route then follows a right-turning at the next intersection and enters Huocheng Road, a narrow one-lane-per-direction road where trucks from the opposite direction sometimes encroaching on the ego lane, necessitating slight steering to avoid collisions. Proceeding through the next intersection, the route goes straight and executes an unprotected left turn to cross a bustling transportation route frequently used by trucks. Finally, the route circles back to Yecheng Road and returns to destination No.1688.
A.11 Model structure
Concurrently, the vectorization branch processes parametric instance information through a PointNet subgraph, which includes three edge convolutions. The output from this branch is then passed through a downstream Transformer encoder to obtain the instance encoding.
Following the extraction of the scene embedding, features by oriented crop around the agents are obtained. These features are then processed through a sequence of maxpooling and a multi-layer perceptron (MLP). The resulting features are concatenated with the instance encoding to create the agent encoding, which is subsequently decoded by an MLP to obtain each agent’s trajectory.
A.12 Training details
The Adam optimizer is used with an initial learning rate of $5xe^{-4}$ and a batch size of 32. The scheduler of reduce on loss plateau decay is also implemented, with a reducing factor of 0.3 and a patience factor of 2, which means that if there is no improvement after 2 epochs, the learning rate will be reduced by 0.3 times the original value. The lower bound of the learning rate is set to $10^{-8}$ and the model parameters are randomly initialized. Experiments are conducted with 4 NVIDIA H800 GPUs with 60 epochs to complete.
Add two stage training @weijian.sun

### v2

@@ A.7 Lattice Sampler
A.8 Retrieval-based Sampler
The retrieval-based sampler does not depend on the map or lane information to generate trajectories, instead, it retrieves a cluster of trajectories whose initial state are close to the ADV’s current state from a pre-established expert trajectory dataset.
@@ A.8 Retrieval-based Sampler
To subdivide the original trajectories into different bins, we use the initial velocity v, acceleration a and curvature $\kappa$, with respective bin sizes of $1.0\frac{m}{s}$, $0.5\frac{m}{s^{2}}$, $0.01\frac{1}{m}$. In each bin, trajectories are clustered into 2000 sets, and the closest trajectory to each cluster center is retained. Finally, the trajectories from each subset will be aggregated to the comprehensive expert trajectory dataset.
@@
$\textit{D}=\beta_{1}\cdot|v-v_{ADV}|+\beta_{2}\cdot|a-a_{ADV}|+\beta_{3}\cdot|\kappa-\kappa_{ADV}|$ (1)
@@ A.9 Hardware
A.10 Task Perf
Figure 4: Tasks on board
Task Time Cost(ms)
Ave 90p Max
Build Frame 0.77 1.13 5.47
Build Obstacles 0.37 0.67 5.35
Build Reference Line 0.52 0.82 10.61
BackBone / Imitation planner 40.90 51.05 81.43
Curve Based Sampler 5.12 10.75 15.64
Retrival Based Sampler 8.31 12.16 19.22
Lattice Sampler 9.83 15.15 40.51
Cost Evaluator 0.58 1.39 9.95
Safety Layer 0.32 0.53 8.12
Over All 50.24 72.26 108.72
Table 4: Task perf on board.
The perception module is implemented on the master chip, whereas the GAD resides on the subsequent chip. We enumerate the time expenditures associated with each task in Tab. 4. These tasks are orchestrated by a flow-based scheduler, facilitating concurrent execution as depicted in Fig. 4. The deployment of the model via TensorRT is conducted without the utilization of acceleration techniques such as quantization or pruning.
@@ A.10 Task Perf
A.11 Road Test
Our road tests are conducted in the Jiading District of Shanghai, China, known as ”Shanghai Automobile City”. The detailed testing route is illustrated in Figure 5, encompassing straight stretches, lane changes, intersections, and turning scenes involving various traffic participants.
The route begins at No.1688 Yecheng Road, a road with two lanes in each direction that encounters frequent cars and electric bicycles. It passes through two intersections controlled by traffic lights before entering a three-lane structure and making a careful lane change to the right-most lane, where sometimes occupied by static vans. It then queues before the stop line and then turns onto Shengxin Road, a four-lane-per-direction road heavily congested commuter route with significant traffic flows in all directions.
@@ A.11 Road Test
The route then follows a right-turning at the next intersection and enters Huocheng Road, a narrow one-lane-per-direction road where trucks from the opposite direction sometimes encroaching on the ego lane, necessitating slight steering to avoid collisions. Proceeding through the next intersection, the route goes straight and executes an unprotected left turn to cross a bustling transportation route frequently used by trucks. Finally, the route circles back to Yecheng Road and returns to destination No.1688. The total length of the test route is 3.0 kilometers long.
A.12 Dateset
We collected the record files in 15 hours, which captured the perception output, the ADV trajectory, and route info. All the records are post-processed to obtain 200k clips in total. By default, we take the 1.5s of past context and predict/plan the future 3.0s contexts, corresponding to 15 frames in the past and 30 frames in the future. We partition the dataset into 16w for training and 4w for testing.
We collected 15 hours of record files, containing messages such as perception outputs, ADV trajectories, traffic lights and route information. Following the collection, a thorough data cleaning process was implemented to analyze and identify the recorded data. Errors such as unconverged ADV localizations and perception frames of irregular frequencies were identified and excluded in order to enhance the quality and reliability of the dataset. The records were then categorized into different scenarios including cruising, lane-change, and turning scenarios. Subsequently, less-common but crucial scenarios, such as unprotected left turning scenarios, were amplified by data augmentation, in order to achieve a well-balanced dataset. While data augmentation resulted in higher loss during training since difficult scenarios were added to the dataset, it enhanced the model’s performance and generalization capability. Finally, all the records were post-processed to obtain 200k clips in total. By default, we take the 1.5 seconds of past contexts and plan for the future 3.0 seconds of contexts, corresponding to 15 frames in the past and 30 frames in the future. We partition the dataset into 160k for training and 40k for testing.
A.13 Training details
The Adam optimizer is used with an initial learning rate of $5e^{-4}$ and a batch size of 32. The scheduler of reduce on loss plateau decay is also implemented, with a reducing factor of 0.3 and a patience factor of 2, which means that if there is no improvement after 2 epochs, the learning rate will be reduced by 0.3 times the original value. The lower bound of the learning rate is set to $10^{-8}$ and the model parameters are randomly initialized. Experiments are conducted with 4 NVIDIA A800 GPUs with 60 epochs to complete.
In the first stage of our proposed two-stage training model, we regress a grid map by a cnn encoder-decoder backbone, which represents the occupation that vehicles may distribute in the future. Specifically, cnn-encoder embeds the scene image into feature map with 12 convolution layers, and then 4 deconvolution layers are used to decode feature map to the output map.
In the second stage of obstacle trajectory prediction, we reuse the CNN encoding layer from the first stage and initialize the encoder parameters with the last training step. As for obstacle trajectory prediction, we use an LSTM layer to replace the cnn decoder. In order to concentrate more on obstacles, we crop the encoding feature map centered agents, as the local scene feature of obstacles. The cropping direction aligns with the orientation of the obstacles. The local scene feature map will be embedding as a local scene vector through a Multi-Layer Perceptron layer and a max-pooling layer. Then, the local scene vector will be sent to the LSTM decoder, concatenated with agent’s history vectorized embedding vector.